# Quantifying Multilingual Performance of Large Language Models Across Languages

### v2

@@ Appendix A Appendix
A.1 Ranking Result For LLMs
@@
Figure 6: Simlarity scores curves of LlaMa-13b and Qwen-13b.
Figure 7: Simlarity scores curves of Bloom-3b and Bloom-7b.
A.2 Details of experiment in Table 4 and Table 3
We selected data from the Tatoeba-Challenge repository [^2] https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/data. Since the number of samples for some low-resource language pairs is small, we extracted 100 samples for each language pair. If there are less than 100, we extracted all samples and extracted them according to the test-dev-train priority.
@@ A.2 Details of experiment in Table 4 and Table 3
Table 10: Similarity score of different language pairs of mistral-7b.
@@
Table 11: Similarity score of different language pairs of qwen-7b.
A.3 Performance of LLMs with Other Sizes
A.3.1 LlaMa-13b and Qwen-13b
We also explore the performance of the similarity metric in LLM of 13B parameters, We use LlaMa-13b and Qwen-13b to display the results. From Figure 6 we can observe that the partial order results of LLM-13B are roughly the same as those of LLM-7B, and there is a clear gap between high-resource languages and low-resource languages.
A.3.2 Bloom-3b and Bloom-7b
Figure 7 shows the result of Bloom-3b and Bloom-7b. Except for the last few layers of the model, there are obvious differences between high-resource languages and low-resource languages which are similar to the above LLMs, while there are smaller differences within the same category of languages. The score is higher than LlaMa, Gemma, Mistral, and Qwen.

### v3

@@ Appendix A Appendix
Ranking Result For LLMs
@@
Figure 6: Simlarity scores curves of Bloom 3B and Bloom 7B.
Details of experiment in Table 4 and Table 3
We selected data from the Tatoeba-Challenge repository [^3] https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/data. Since the number of samples for some low-resource language pairs is small, we extracted 100 samples for each language pair. If there are less than 100, we extracted all samples and extracted them according to the test-dev-train priority. We can observe that the scores of language pairs of the same type may vary differently but the performance of each LLM on different types of language pairs is roughly the same.
@@
Table 10: Similarity score of different language pairs of mistral 7B.
@@
Table 11: Similarity score of different language pairs of qwen 7B.
Performance of LLMs with Other Sizes
Bloom 3B and Bloom 7B
Figure 6 shows the result of Bloom 3B and Bloom 7B. Except for the last few layers of the model, there are obvious differences between high-resource languages and low-resource languages which are similar to the above LLMs, while there are smaller differences within the same category of languages. The score is higher than LlaMa2, Gemma, Mistral, and Qwen.
The Relation Between the Performance of Reasoning and The Similarity Score
We also quantify the relationship between the performance of the inference task (Table 2) and the language similarity to English (Figure 1). It can be clearly observed from Figure 7 and Figure 8 that the performance of LLM in reasoning tasks is highly correlated with the similarity score, it performs better in high-resource languages(the first four languages) and worse in low-resource languages(the last four languages).
Figure 7: Accuracy of ARC reasoning tasks and similarity scores in different languages on Gemma 7B.
Figure 8: Accuracy of ARC reasoning tasks and similarity scores in different languages on Mistral 7B.